# New Recipe for Semi-supervised Community Detection: Clique Annealing under Crystallization Kinetics

### v1

@@ 3 Problem Definition and Pipeline
Problem Definition: Give a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X})$, where the $\mathcal{V}$ represent the node set, $\mathcal{E}$ represents the edge set and $\mathcal{X}$ represents the node feature. The expert-labeled training communities are denoted as $C=\{C_{1},...,C_{N}\}$ where $N$ is the number of expert-labeled communities. The objective in semi-supervised community detection, is formulated as given a small set of expert-labeled communities $C^{train}=\{C_{1},...,C_{m}\}$ as training data, where $m$ ($m\ll N$), to predict $N$-$m$ communities $C^{pred}$ from $\mathcal{G}$. The predicted communities should be consistent with all rest labeled communities (test communities), $C^{test}$ = $C\setminus C^{train}$.
@@ 4.2 Implementation of Crystallization Kinetics
$\bar{\circ}$ Replace mnodes in $\circ$
$\dot{\circ}$ Remove mnodes from $\circ$
$\breve{\circ}$ one-hop neighbors of $\circ$
@@
$loss^{E}_{Size}=\sum_{(i,j)}^{Pos-S}\mbox{max}\{0,||i||-||j||\}+\sum_{(i,j)}^{Neg-S}\mbox{max}\{0,\alpha-(||i||-||j||)\},\$ (1)
where $\alpha$ is the loss margin, and |||||$\cdot$||| | ⋅ | | denotes the norm of the subgraph embedding $h$, which we utilize as an indicator of the graph’s stored energy. We further define the positive and negative pair $Pos$-$S$ = $\{(b,a_{1}),(c,a_{1}),(\dot{a_{1}},a_{1})\}$, $Neg$-$S$ = $\{(a_{1}+a_{2},a_{1}),(a_{2}+a_{3},a_{2}),(a_{3}+a_{1},a_{3})\}$, where ‘$+$’ indicates combination. For each pair in $Pos$-$S$, the stored energy of the first subgraph is less than the second’s, reflecting smaller subgraph sizes. Conversely, in $Neg$-$S$, the first subgraph, representing a merged subgraph, exhibits greater stored energy than the second subgraph due to its bigger size.
@@
$loss^{E}_{Defc}=\sum_{(i,j)}^{Pos-D}\mbox{max}\{0,||i||-||j||\},$ (2)
where $Pos$-$D$ = $\{(a_{1},\bar{a_{1}}),(b,\bar{b}),(c,\bar{c})\}$. For each pair in $Pos$-$D$, the latter is distorted from the first subgraph (inside nodes are replaced with outside connected nodes). The second subgraph’s stored energy should thus be larger than the first one.
@@
$loss^{I}=\sum_{(i,j)}^{Pos-I}\max\{0,||i||-||j||-\sum_{v\in{\breve{j}}}\mbox{Intf-E}(v)\}+\sum_{(i,j)}^{Neg-I}\max\{0,||i||-||j||+\sum_{v\in{\breve{i}}}\mbox{Intf-E}(v)\},$ (4)
@@
where $W_{I}$ and $b_{I}$ represent the weights and bias respectively, $f_{v}=[f_{v}^{a}||f_{v,g}^{e}]$, $f_{v}^{a}$ = [$f_{v}^{o}$, $deg(v)$, $max(DN(v))$, $min(DN(v))$, $avg(DN(v))$, $std(DN(v))$] is the standard augment feature of node $v$, $f_{v}^{o}$ is the raw features of node $v$ with default value of 1, following the previous studies  (Wu et al., 2022; Cai & Wang, 2018; Zhang et al., 2020). $deg(v)$ is the degree of node $v$ and $DN(v)$ represents the degree set of the neighbor nodes of $v$.
For the corresponding subgraph $g$, the external feature $f_{v,g}^{e}$ is represented as [$l$-$num$, $graph$-$size$, $mis$-$num$], where $l$-$num$ is the edge number between node $v$ and $g$. $graph$-$size$ is the node number of $g$. It is worth noting that an ideal community is distinguished by high exclusivity and strong internal connections, forming a clique. Accordingly, $mis$-$num$ is defined as the number of nodes that are not part of a clique.
@@
where $\sigma$ stands for activation function, ||||| | stands for concatenation, $h(*)$ and $\bar{h}(*)$ are the graph embedding and normalized embedding for the given subgraph. $W_{p}^{k}$ and $b_{p}^{k}$ are the weights and bias of the corresponding network. The integrity-based loss function can thus be formed as:
$loss^{G}=-\frac{1}{3}\sum_{S}^{\{S_{1},S_{2},S_{3}\}}\sum_{g\in S}\sum_{k=1}^{3}({y_{g}^{k}}\mbox{log}(\hat{y}_{g}^{k})+(1-y_{g}^{k})\mbox{log}(1-\hat{y}_{g}^{k})),\$ (7)
@@ 4.3 Graph Encoder
$\mathscr{L}=\gamma^{E}loss^{E}+\gamma^{C}loss^{C}+\gamma^{I}loss^{I}\$ (9)
@@ 5.1 Potential for Further Growth of the Current State
||*|| [E] Stored energy of a subgraph
@@ 5.2 Selection of Clique for Merging
$\hat{S}_{i}^{c}$ =$=\{S_{i}^{c}|P_{i}>P_{temp}(|S_{i}^{c}|)\},\;(i\in\{1,\dots,|V^{e}|\}),$ (13)
$S_{m}$ =$=\bigcup\{S_{i}^{c}|\mathbbm{1}(C^{e}_{i},S_{m-1},S_{i}^{c})=1\},\;S_{i}^{c}\in{\hat{S}_{i}^{c}},$
@@ 6.1 Overall Performance
Table 3: F1 Scores (Jacc., OMNI in Table 9). A/D: find A’s communities in A+D. Bold/Underline: 1st/2nd best scores. N/A: not converge in $2$ days. We conduct $5$ experiments for NP and CLANN, the average std is less than $0.0323$.
@@ Appendix A Clique-Based Modularity and Loss in Community Detection
Energy-Based Loss: To understand why the modularity of a clique with size $k$ is always greater than that of a sub-clique with size $k-1$, consider the modularity formula in detail. When you remove a node from a $k$-clique to form a $(k-1)$-clique, you also remove all the edges connected to that node. This reduces the number of positive terms in the modularity summation. Specifically, for each node pair $(i,j)$ within the original clique, the term $\left(a_{i,j}-\frac{k_{i}\cdot k_{j}}{2m}\right)$ contributes positively to the modularity when $i$ and $j$ are connected (which is always true in a clique). By removing a node from the community, you reduce the number of such positive contributions, thereby lowering the overall modularity. Therefore, the modularity of the original $k$-clique is always greater than that of any $(k-1)$-clique derived from it.
@@ Appendix D Preliminaries of Hyperbolic Geometry
Hyperbolic geometry encompasses several conformal models Cannon et al. (1997). Based on its widespread use in deep learning and computer vision, we operate on the Poincaré ball. The Poincaré ball is defined as ($\mathbb{D}_{c}^{n}$, $g^{\mathbb{D}_{c}^{n}}$), with manifold $\mathbb{D}_{c}^{n}$ = $\{x\in\mathbb{R}^{n}:c||x||<1\}$ and Riemannian metric:
$g^{\mathbb{D}_{c}}_{x}=(\lambda_{x}^{c})^{2}g^{E}=\frac{2}{1-c{||x||}^{2}}\mathbb{I}^{n},$ (14)
@@
$\mbox{exp}_{v}^{c}(x)$ $=v\oplus_{c}(\mbox{tanh}(\sqrt{c}\frac{\lambda_{v}^{c}||x||}{2})\frac{x}{\sqrt{c}||x||}),$ (15)
$v\oplus_{c}w$ $=\frac{(1+2c\langle v,w\rangle+c{||w||}^{2})v+(1-c{||v||}^{2})w}{1+2c\langle v,w\rangle+c^{2}{||v||}^{2}{||w||}^{2}}.$
@@ Appendix E Dataset Statistics
Table 7: Dataset statistics of community number #$\hat{C}$, vertex number #$V$, edge number #$E$, max(average) community size $C_{M/A}$, the logarithm of edge number/vertex number $log(E/V)$, and coverage ratio $R_{c}$. $A$, $D$, and $L$ stand for Amazon, DBLP, and LiveJournal, respectively. Additionally, Est.$\alpha$ represents the estimated $\alpha$ value of degree power law fit.
#$\hat{C}$V #E $C_{M/A}$ log(E/V) $R_{c}$ Est.$\alpha$
@@
Table 8: Dataset statistics of different community numbers (1k-5k). The meanings of each notation are identical to the previous table. Additionally, Est.$\alpha$ represents the estimated $\alpha$ value of degree power law fit.
#V #E $C_{M/A}$ log(E/V) $R_{c}$ Est.$\alpha$
@@ Appendix F Preliminary Core Filter
The Nucleus Proposer needs to prepare all nodes’ cliques. However, for large graphs, finding all cliques is prohibitively time-consuming. To address this challenge, we develop a preliminary selection mechanism for large graphs before the Nucleus Proposer. Concretely, for each community, nodes with the highest betweenness centrality are labeled as community cores. The remaining nodes are considered peripheral nodes and labeled as 00. To characterize the feature of node $v$, we concatenate the original representation $f_{v}^{a}$ and the average original features $\bar{f}_{v}^{a}$ of $v$’s neighbors. We train the preliminary classifier with communities from the training set:
@@ Appendix G Evaluation Metrics & Baseline
[…]
@@ Appendix I Supplementary Parameters Analysis
Dataset ${Engy.}$ +${Intf.}$ +${Cons.}$ +${Intg.}$ NP +Infc +SA +C-E +TA
Jaccard A 0.6925 0.7006 0.7228 0.7631 0.7227 0.7458 0.7696 0.8148 0.8600
@@
Est.$\alpha$ SBM N-SBM O-SBM Louvain Label Prop. Spectral CLANN

### v2

@@ 3 Problem Definition and Pipeline
Problem Definition: Give a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X})$, where the $\mathcal{V}$ represent the node set, $\mathcal{E}$ represents the edge set and $\mathcal{X}$ represents the node feature. The expert-labeled training communities are denoted as $C=\{C_{1},...,C_{N}\}$ where $N$ is the number of expert-labeled communities. The objective in semi-supervised community detection, is formulated as given a small set of expert-labeled communities $C^{train}=\{C_{1},...,C_{m}\}$ as training data, where $m$ ($m\ll N$), to predict NN-$m$ communities $C^{pred}$ from $\mathcal{G}$. The predicted communities should be consistent with all rest labeled communities (test communities), $^{test} = CC$$\setminus C^{train}$.
@@ 4.2 Implementation of Crystallization Kinetics
$\bar{\circ}$ Replace m% nodes in $\circ$
[…]
@@
[…]
where $\alpha$ is the loss margin, and $||\cdot||$ denotes the norm of the subgraph embedding $h$, which we utilize as an indicator of the graph’s stored energy. We further define the positive and negative pair PosPos-$S$ = $\{(b,a_{1}),(c,a_{1}),(\dot{a_{1}},a_{1})\}$, NegNeg-$S$ = $\{(a_{1}+a_{2},a_{1}),(a_{2}+a_{3},a_{2}),(a_{3}+a_{1},a_{3})\}$, where ‘++’ indicates combination. For each pair in PosPos-$S$, the stored energy of the first subgraph is less than the second’s, reflecting smaller subgraph sizes. Conversely, in NegNeg-$S$, the first subgraph, representing a merged subgraph, exhibits greater stored energy than the second subgraph due to its bigger size.
@@
[…]
where PosPos-$D$ = $\{(a_{1},\bar{a_{1}}),(b,\bar{b}),(c,\bar{c})\}$. For each pair in PosPos-$D$, the latter is distorted from the first subgraph (inside nodes are replaced with outside connected nodes). The second subgraph’s stored energy should thus be larger than the first one.
@@
$\mbox{Intf-E}(v)=\mbox{Softplus}(W_{I}*f_{v}+b_{I}),$ (5)
where $W_{I}$ and $b_{I}$ represent the weights and bias respectively, $f_{v}=[f_{v}^{a}||f_{v,g}^{e}]$, $f_{v}^{a}$ = $f_{v}^{o}$, $deg(v)$, $max(DN(v))$, $min(DN(v))$, $avg(DN(v))$, $std(DN(v))$] is the standard augment feature of node $v$, $f_{v}^{o}$ is the raw features of node $v$ with default value of 1, following the previous studies  (Wu et al., 2022; Cai & Wang, 2018; Zhang et al., 2020). $deg(v)$ is the degree of node $v$ and $DN(v)$ represents the degree set of the neighbor nodes of $v$.
For the corresponding subgraph $g$, the external feature $f_{v,g}^{e}$ is represented as [$l$-numnum, graphgraph-sizesize, mismis-numnum], where $l$-numnum is the edge number between node $v$ and $g$. graphgraph-sizesize is the node number of $g$. It is worth noting that an ideal community is distinguished by high exclusivity and strong internal connections, forming a clique. Accordingly, mismis-numnum is defined as the number of nodes that are not part of a clique.
@@
where $\sigma$ stands for activation function, $||$ stands for concatenation, h()h(*) and $\bar{h}(*)$ are the graph embedding and normalized embedding for the given subgraph. $W_{p}^{k}$ and $b_{p}^{k}$ are the weights and bias of the corresponding network. The integrity-based loss function can thus be formed as:
$loss^{G}=-\frac{1}{3}\sum_{S}^{\{S_{1},S_{2},S_{3}\}}\sum_{g\in S}\sum_{k=1}^{3}({y_{g}^{k}}\mbox{log}(\hat{y}_{g}^{k})+(1-y_{g}^{k})\mbox{log}(1-\hat{y}_{g}^{k})),\\$ (7)
@@ 4.3 Graph Encoder
$\mathscr{L}=\gamma^{E}loss^{E}+\gamma^{C}loss^{C}+\gamma^{I}loss^{I}\\$ (9)
@@ 5.1 Potential for Further Growth of the Current State
||*$|$ [E] Stored energy of a subgraph
@@
Interface Energy Check. As mentioned in Sec. 4.2, to surpass the interface barrier, subgraphs need to consume extra energy. For an extendable node $v^{e}_{i}$ $\in$ $V^{e}$, we merge all cliques containing $v^{e}_{i}$ as $C^{e}_{i}$, the corresponding expanded candidate $S_{i}^{c}$ is constructed by combining $C^{e}_{i}$ with $S_{m-1}$. Due to the interface energy barrier, the stored energy summation of the current state $S_{m-1}$, merged clique $C^{e}_{i}$, and their interface energy $\mbox{Intf-E}(v^{e}_{i})$ (defined in Eq. 5) should be larger than the stored energy of the expanded candidate. We therefore require the following interface energy constraint:
@@ 5.2 Selection of Clique for Merging
$\hat{S}_{i}^{c}$ $=\{S_{i}^{c}|P_{i}>P_{temp}(|S_{i}^{c}|)\},\;(i\in\{1,\dots,|V^{e}|\}),$ (13)
$S_{m}$ $=\bigcup\{S_{i}^{c}|\mathbbm{1}(C^{e}_{i},S_{m-1},S_{i}^{c})=1\},\;S_{i}^{c}\in{\hat{S}_{i}^{c}},$
@@ 6.1 Overall Performance
Table 3: F1 Scores (Jacc., OMNI in Table 9). A/D: find A’s communities in A+D. Bold/Underline: 1st/2nd best scores. N/A: not converge in 22 days. We conduct $5$ experiments for NP and CLANN, the average std is less than $0.0323$.
@@ 7 Conclusion and Future Work
8 Acknowledgement
This research was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (Project ID: 22-SIS-SMU-048). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
@@ Appendix A Clique-Based Modularity and Loss in Community Detection
Energy-Based Loss: To understand why the modularity of a clique with size $k$ is always greater than that of a sub-clique with size k1k-1, consider the modularity formula in detail. When you remove a node from a $k$-clique to form a (k1)(k-1)-clique, you also remove all the edges connected to that node. This reduces the number of positive terms in the modularity summation. Specifically, for each node pair $(i,j)$ within the original clique, the term $\left(a_{i,j}-\frac{k_{i}\cdot k_{j}}{2m}\right)$ contributes positively to the modularity when $i$ and $j$ are connected (which is always true in a clique). By removing a node from the community, you reduce the number of such positive contributions, thereby lowering the overall modularity. Therefore, the modularity of the original $k$-clique is always greater than that of any (k1)(k-1)-clique derived from it.
@@ Appendix D Preliminaries of Hyperbolic Geometry
[…]
@@ Appendix E Dataset Statistics
[…]
@@
[…]
@@ Appendix F Preliminary Core Filter
The Nucleus Proposer needs to prepare all nodes’ cliques. However, for large graphs, finding all cliques is prohibitively time-consuming. To address this challenge, we develop a preliminary selection mechanism for large graphs before the Nucleus Proposer. Concretely, for each community, nodes with the highest betweenness centrality are labeled as community cores. The remaining nodes are considered peripheral nodes and labeled as 0. To characterize the feature of node $v$, we concatenate the original representation $f_{v}^{a}$ and the average original features $\bar{f}_{v}^{a}$ of $v$’s neighbors. We train the preliminary classifier with communities from the training set:
@@ Appendix G Evaluation Metrics & Baseline
$\frac{1}{2}(\frac{1}{N}\sum_{i}\underset{j}{\mathrm{max}}\delta(\hat{C}^{i},\dot{C}^{j})+\frac{1}{M}\sum_{j}\underset{i}{\mathrm{max}}\\ \delta(\dot{C}^{j},\hat{C}^{i})),$ (19)
@@ Appendix I Supplementary Parameters Analysis
Dataset Engy.{Engy.} +Intf.{Intf.} +Cons.{Cons.} +Intg.{Intg.} NP +Infc +SA +C-E +TA
Jaccard
A 0.6925 0.7006 0.7228 0.7631 0.7227 0.7458 0.7696 0.8148 0.8600
@@
[…]